a long distance trucker traveled 72 miles in one direction during a snow storm. The return trip in rainy weather was accomplished at double the speed and took 2 hours less time. find the speed going
Answer by mananth(16946)

a long distance trucker traveled 72 miles in one direction during a snow storm. The return trip in rainy weather was accomplished at double the speed and took 2 hours less time. find the speed going
speed while going = x mph
distance traveled = 72 miles
..
speed while returning = 2x
distance = 72 miles.
..
Time going - time return = 2 hours
..
t=d/r
72/x - 72/2x= 2
LCD=2x
(144-72)/2x=2
72=4x
/4
x=18 mph ====> speed while going